# MobileLLM-R1: Exploring the Limits of Sub-Billion Language Model Reasoners with Open Training Recipes

### v2

@@ 4 Final Results
Post-trained Model Table 7 presents the performance of post-trained models. Notably, on LiveCodeBench, small models below 400M parameters struggle to produce reliable outputs. In contrast, MobileLLM-R1-360M achieves 5.1 points, surpassing even models with over 1B parameters, such as SmolLM2-1.7B, Gemma3-1B, and LLaMA3.2-1B. Remarkably, MobileLLM-R1-950M demonstrates a substantial accuracy gain over Qwen3-0.6B on LiveCodeBench and even matches the performance of much larger state-of-the-art models, such as DeepSeek-R1-Distill-Qwen-1.5B. Across Math and AIME benchmarks, MobileLLM-R1 consistently outperforms other fully open-source models and achieves scores comparable to the partially open-source Qwen3 series.
5 Pre-training: Balance of Capabilities
The notion of reasoning in large language models (LLMs) remains both complex and contested. While the term is sometimes used to describe a model’s ability to engage in structured, multi-step inference (Wei et al., 2022; Kojima et al., 2022), it has also become a proxy for improved performance on challenging benchmarks (Srivastava et al., 2022). In this work, we adopt a pragmatic stance: we treat gains on reasoning-centric benchmarks as reasonable evidence of enhanced reasoning behaviors, while remaining cautious about equating such gains with genuine reasoning ability in the cognitive sense (Bender & Koller, 2020). Concretely, this entails selecting informative datasets that most effectively enhance the target capability (Section 5.1) and optimizing their combination ratios to maximize knowledge acquisition within the fixed token budget (Section 5.2). Figure 4 illustrates the training pipeline with the full procedure in Appendix 3.
5.1 Selecting Informative Datasets for Target Capability
Hierarchical Rejection Sampling. To derive a compact capability-probing datasets for each domain, we employ a hierarchical rejection sampling pipeline that integrates multiple classifier- and model-based filters, as shown in Figure 5 The objective is to construct a small yet representative target dataset for each capability, such that it can serve as a faithful proxy for reasoning performance while dramatically reducing overall volume during evaluation. For each corpus in Table 3, we first apply the FineWeb-Edu classifier (Penedo et al., 2024) to select samples with high educational value, retaining only those with classifier scores above $4$. Next, we incorporate model-based evaluation by scoring each remaining sample using the Ask-LLM paradigm (Sachdeva et al., 2024). The evaluation prompt asks the model to judge whether a sample should be included in a reasoning-probing dataset, framed as a binary classification task (“1” for inclusion, “0” for exclusion). Rather than relying solely on the hard prediction, we record the probability assigned to “1” as a graded measure of the model’s confidence in the example’s reasoning relevance. For all Ask-LLM scoring, we select the top 10% samples within each dataset. This step complements classifier-based quality filtering by directly capturing signals of reasoning relevance, consistent with recent findings that costly, fine-grained quality samplers can outperform simple maximum-coverage approaches in terms of data efficiency (Sachdeva et al., 2024; Pang et al., 2025; Chen & Zhou, 2025). Next, we apply a domain-specific prompt to Ask-LLM for each capability with specific emphasis on code, math, general knowledge or combined. Finally, we perform semantic deduplication across corpora, shrinking each dataset in Table 3 to a subset of roughly 10,00010{,}000 examples. This yields the representative datasets $\mathcal{D^{R}}_{i}$, each containing highly representative samples for its corresponding corpus.
Figure 5: Hierarchical rejection sampling. We employ the FineWeb-Edu classifier in conjunction with the ASK-LLM paradigm to construct a representative subset from each pretraining corpus. Samples are assigned selection scores, and inclusion is determined by thresholding these scores.
@@ 5.1.1 Curation of Representative Datasets and Capability-Probing Dataset
5.1.2 Disentangling the Impact of Data Sources
[…]
Leave-One-Out Ablations. We operationalize Eq. 1 by training models under leave-one-out settings and measuring the resulting differences in loss across benchmarks. Together, these analyses highlight not only which sources matter most, but also how much marginal benefit additional data from a given source provides. This methodology allows us to disentangle the contributions of heterogeneous data sources to reasoning-related performance in code, knowledge, and mathematics.
In contrast, domain-specific datasets primarily strengthen their respective domains: Starcoder substantially improves code performance (and, interestingly, math), while math-focused corpora primarily benefit math. However, their transfer to general knowledge is limited. An unexpected observation is that Starcoder benefits math more than OpenWeb-Math benefits code, a reversal of the commonly held view that mathematical data contributes disproportionately to coding ability (Lewkowycz et al., 2022). Finally, Wikipedia appears to contribute little to math or code compared to web or domain-specific data, yet remains necessary as a structured and reliable source of factual knowledge.
5.2 Datamixing via Cross-Capability Self-Influence
$\text{Self-influence: }\mathcal{I}(x_{i},x_{test}\in\mathcal{D^{P}_{C}};\theta_{\mathcal{C},t}),$ (3)
[…]
Then, the joint influence of a sample is computed as
Each source dataset $g$ is then assigned a sampling weight ($w_{g}$):
6 Mid-training: Knowledge Compression
[…]
(2) Dataset-level influence for adaptive data mixing. Beyond sample-level filtering, we aggregate influence scores to the dataset level, enabling adaptive control of the mixing ratio among mid-training datasets, according to Eqs. 5 and 6).
$\theta_{t+1}=\mathrm{MidTrain}(\theta_{t},\mathcal{D}_{t}).$ (8)
Intuition: Distributional Compression of Influence. The compression phases can be viewed as iteratively distilling the mid-training dataset in alignment with the model’s evolving capacity throughout training. In early phases, the influence scores are more varied because the model $\theta_{t}$ is still under-trained. However, as $t$ increases, the model becomes better aligned with the target distribution, and its estimates of sample importance are narrowed down (See Figure 8). This recursive interplay produces increasingly refined datasets: uninformative (or negative-influence) samples are discarded, thus amplifying the impact of high-value samples. Conceptually, compression phases mimic an iterative denoising process, where each step sharpens the signal from $\mathcal{D}^{(\mathrm{target})}$ against the noisy background of $\mathcal{D}^{(\mathrm{raw})}$. We terminate the iteration until the distribution of influence converges to approximately zero.
7 Post-Training
Using the datasets from Sections 5 and 6, we obtain MobileLLM-R1-base. Given that our primary goal is to elucidate how data curation in pre-training and mid-training builds strong small reasoning models, we leverage established supervised fine-tuning (SFT) datasets. We first apply Tülu-3-SFT (Lambert et al., 2024) dataset for instruction alignment and OpenScienceReasoning-2, OpenCodeReasoning-2 (Ahmad et al., 2025) and OpenMathReasoning (Moshkov et al., 2025) for reasoning-oriented SFT to extend context and elicit long chain-of-thought reasoning.
7.1 Post-training data
Tülu-3-SFT (Lambert et al., 2024) is a large, curated collection of prompts and instruction‑answer pairs designed for post-training language models across skills such as reasoning, coding, and math. It combines publicly available datasets with synthetically generated data. We use tulu-3-sft-olmo-2-mixture-0225, which is a latest filtered version of Tulu-3 dataset.
OpenMathReasoning (Moshkov et al., 2025) is a large-scale dataset for training models in complex mathematical reasoning. It contains 3.2M long chain-of-thought (CoT) solutions for 306K unique mathematical problems, helping models improve problem-solving skills in mathematics.
OpenScienceReasoning-2 (NVIDIA, 2025) is a synthetic dataset spanning multiple scientific domains. It contains 803k general-purpose reasoning data, including STEM, law, economics, and humanities, featuring both multiple-choice and open-ended questions, designed to enhance scientific reasoning capabilities.
OpenCodeReasoning-2 (Ahmad et al., 2025) is a large-scale dataset for programming reasoning. It contains 1.4M Python and 1.1M C++ samples, covering nearly 35K unique competitive programming problems. It is intended to improve code completion abilities in models.
7.2 Post-training ablation
Stage 1 Stage 2 MATH GSM8K LCBv6 MMLU
Ablation on Stage 1
w/o Tulu-3 M + C + S 56.2 68.2 13.1 44.0
w/ Tulu-3 M + C + S 57.8 68.5 13.7 43.7
Ablation on Stage 2
Tulu-3 M 57.4 68.2 0.0 43.1
Tulu-3 C 16.2 31.0 12.0 39.9
Tulu-3 S 23.8 62.2 3.4 45.6
Tulu-3 M + C 58.4 65.6 10.9 40.4
Tulu-3 M + S 60.0 66.9 0.6 45.0
Tulu-3 C + S 29.4 65.3 14.3 44.4
Tulu-3 M + C + S 57.8 68.5 13.7 43.7
Joint Ablation
Tulu-3 + (M + C + S) 56.2 53.1 14.9 44.0
Tulu-3 (M + C + S) 57.8 68.5 13.7 44.0
Table 9: Evaluation of reasoning capabilities elicited by different models when fine-tuned on the same reasoning supervised finetuning (SFT) dataset. Baseline models use their instruct checkpoints; OLMo uses its SFT checkpoint before reinforcement learning; our model uses intermediate Tulu3-SFT checkpoints, denoted with *. All models are trained for one epoch on the joint reasoning SFT corpus (OpenMathReasoning + OpenScienceReasoning-2 + OpenCodeReasoning-2).
Model Size MATH GSM8K LCBv6
SmolLM2-135M-Instruct 135M 3.2 1.6 0.6
MobileLLM-R1-140M* 140M 4.8 3.7 1.1
SmolLM2-360M-Instruct 362M 5.2 7.4 3.4
MobileLLM-R1-360M* 359M 19.2 23.8 4.0
OLMo-2-0425-1B-SFT 1.48B 53.0 58.8 11.4
SmolLM2-1.7B-Instruct 1.71B 41.4 50.5 7.4
MobileLLM-R1-950M* 949M 57.8 68.5 13.7
We conduct ablation studies to understand the contribution of different supervision signals and training strategies in our two-stage post-training pipeline. Results in Table 8 reflect several interesting insights.
(1) Instruction-following supervision provides a strong foundation for reasoning. Models that are first trained with Tulu-SFT data consistently outperform those that start directly from reasoning data. This demonstrates that high-quality instruction-following supervision provides alignment signals making subsequent reasoning adaptation more effective.
(2) Domain-specific data improves performance in its own domain. OpenMathReasoning, OpenScienceReasoning-2, and OpenCodeReasoning-2 improve performance on math, code, and knowledge-intensive benchmarks (e.g., MMLU), respectively. These effects are highly consistent, showing that specialized datasets provide targeted benefits.
(3) Scientific reasoning data generalizes across domains. Beyond its impact on MMLU, adding science data also improves performance on math and coding tasks compared to using only math and coding data. This suggests that scientific reasoning provides broadly useful structures that transfer across different forms of symbolic problem-solving.
(4) Symbolic Reasoning Gains Come at the Cost of Knowledge Retention. Introducing math or coding data tends to reduce performance on MMLU. We hypothesize this is due to limited model capacity: when the model is pushed to absorb new symbolic reasoning skills through exposure to math and coding data, it partially forgets factual knowledge that is important for knowledge-intensive evaluation. This forgetting effect is especially pronounced in smaller models.
(5) Decoupling Alignment and Reasoning for Stronger Models. Compared to Model trained with Tulu and reasoning data jointly in a single staget, a two-stage setup of first training on Tulu, then on reasoning data delivers overall stronger results especially on math and general reasoning benchmarks.
7.3 Comparison with baselines on identical reasoning SFT
To disentangle the contribution of curated pre-training and mid-training data from that of high-quality post-training data, we conducted an ablation study. Specifically, we finetune all baseline instruct models, as well as the MobileLLM-R1 general supervised fine-tuned model (trained for 2 epochs on the Tulu dataset), on the joint reasoning SFT corpus (OpenMathReasoning + OpenScienceReasoning-2 + OpenCodeReasoning-2) for one epoch. Our results in Table 9 show that, even under identical supervised fine-tuning, models with stronger pre-training and mid-training exhibit more robustly embedded knowledge, which in turn facilitates the elicitation of reasoning capabilities during post-training. In that sense, MobileLLM-R1 consistently outperforms prior models trained on fully open-source corpora, such as OLMo-2 and SmolLM, on reasoning benchmarks. Notably, our 140M and 360M checkpoints achieve substantial gains over SmolLM baselines, while our 950M model surpasses both OLMo-2 1.48B and SmolLM-1.7B, despite its significantly smaller size.
7.4 Discussion: Whether to Use RL or Not
8 On-device Profiling
@@
Table 10 reports the measured generation throughput (tokens per second) of MobileLLM-R1 models and LLaMA-3 baselines across different context lengths on the Samsung Galaxy S22. The results confirm the clear efficiency advantage of smaller models: the 140M variant sustains over 100 tokens/s up to 8k tokens, while even the 360M and 950M models maintain practical throughput across medium-length contexts. In contrast, larger models such as LLaMA-3.2-3B and LLaMA-3.1-8B encounter out-of-memory failures beyond short contexts, and the LLaMA-3.2-1B model runs at less than one-quarter the speed of MobileLLM-R1-140M. Moreover, long-context settings (16k–32k) quickly exhaust memory capacity for billion-scale models, highlighting the prohibitive cost of KV-cache storage at scale. These findings demonstrate that sub-billion parameter models not only enable reasoning with reduced memory overhead but also achieve far higher generation speed, making them far better suited for on-device long-context inference. We benchmark the latency of MobileLLM-R1 models (140M, 360M, 950M) and LLaMA-3 models (1B, 3B, 8B) using ExecuTorch on a Samsung Galaxy S22 (8 GB RAM). In this experiment, we run each model with a warm-up phase and then average the results over three iterations using the same prompt. On-device applications usually adopt quantized settings. To reflect typical on-device deployment, models are benchmarked under quantized settings: linear layers are quantized using 8-bit dynamic activations and 4-bit weights. For the 125M model, a group size of 32 is adopted since its hidden dimension is not divisible by 128, whereas all larger models use a group size of 128. Embedding layers are quantized to 4-bit with a group size of 32.
9 Related Work
The advent of GPT-3 (Brown et al., 2020) demonstrated the transformative potential of large langauage models (LLMs), catalyzing a wave of research on LLMs, including both proprietary models like Claude (Anthropic, 2024), and open-source alternatives such as LLaMA (Touvron et al., 2023a, b; Dubey et al., 2024), Gemma (Team et al., 2024a, b, 2025), Qwen (Team, 2024; Yang et al., 2025), etc. The scaling laws of large language models (LLMs) have long been a central focus in the field. Researchers are particularly interested in understanding how increasing model size can elicit emergent capabilities and whether there exist potential singularities in performance. Concurrently, there is growing attention on the efficiency-accuracy trade-offs achievable with smaller models. Investigating the performance limits of parameter-constrained models not only elucidates the fundamental boundaries of representational capacity but also highlights opportunities for optimizing computational efficiency and resource utilization. Notably, MobileLLM (Liu et al., 2024) represents the first LLM explicitly designed for on-device deployment. Subsequently, numerous high-performance and small-scale models, including OLMO (OLMo et al., 2024), SmolLM (Allal et al., 2025) have been open-sourced. Moreover, these work also denotes a broader trend toward transparency in model development. Beside opensourcing the model weights, they also have provided both weights, training data sources and full training pipelines.
Recently, research focus has gradually shifted from System I to System II thinking, i.e., from instinctive outputs to eliciting explicit reasoning capabilities through methods since the introduction of OpenAI’s O1 model (Jaech et al., 2024). The release of DeepSeek-R1 marked a very successful instance of exposing a model training process explicitly aimed at enhancing reasoning (Guo et al., 2025). Qwen (Yang et al., 2025) represents another strong example of building high-quality reasoning models, with its smaller-scale variants consistently achieving state-of-the-art results across benchmark evaluations.
Prior studies have consistently shown that reasoning capabilities typically emerge only after training on extremely large corpora of tokens (Yang et al., 2025). In contrast, our work shows that a small model can achieve strong reasoning abilities without massive datasets. Table 11 compares state-of-the-art open-source language models under the sub-billion parameter regime, highlighting differences in training scale, context length, openness, and reasoning support. Most prior approaches rely on in-context CoT prompting to elicit reasoning, whereas MobileLLM-R1 and Qwen3 explicitly optimize on long-form reasoning traces, enabling the generation of multi-step reasoning sequences without dependence on prompt-based process. While many recent models, such as Qwen 3 (0.6B, 36T tokens) and LLaMA 3.2 (1.24B, 9T tokens), rely on extremely large proprietary or partially open datasets, only a few—including SmolLM2, OLMO 2, and our proposed MobileLLM-R1 series—provide both weights and fully reproducible training recipes. Notably, despite using significantly fewer training tokens (4.2T) than models like Qwen 3, MobileLLM-R1 achieves competitive or superior reasoning capabilities, underscoring the importance of dataset curation over sheer scale.
Table 11: Comparison of attributes of state-of-the-art sub-billion parameter (<< 1B or $\approx$ 1B) language models
Context OSS OSS data source Model
Name Sizes (# Params) # Training tokens length weights & full recipe type
SmolLM2 135M, 360M 2T (135M), 4T (360M) 8k Instruct
OLMO 2 1B (1.48B) 4.5T 4k Instruct
Qwen 2.5 0.5B 18T 32k Instruct
Qwen 3 0.6B 36T 32k Thinking
LLaMA 3.2 1B (1.24B) 9T 128k Instruct
Gemma 3 270M, 1B 6T (270M), 2T (1B) 32k Instruct
MobileLLM 125M, 350M, 600M, 1B 1T 2k Base
MobileLLM-R1 140M, 360M, 950M 4.2T 32k Thinking
10 Conclusion
Appendix
Appendix A Trajectory of Reasoning Capability Across Training Phases
Appendix B Learning Rate Affects Representation Learning
@@
Therefore, we investigate structural diagnostics of pretrained representations that can act as low-cost proxies for downstream generalization. We draw inspiration from RankMe (Garrido et al., 2023), originally proposed in the context of vision self-supervised finetuning:
$\small\text{RankMe}(Z)=\exp\left(-\sum_{k=1}^{\min(N,K)}p_{k}\log p_{k}\right),\quad\text{where }p_{k}=\frac{\sigma_{k}(Z)}{\sum_{i=1}^{\min(N,K)}\sigma_{i}(Z)}+\epsilon,$ (9)
@@
Appendix C Per-task Accuracy on Common-Sense Reasoning Benchmarks
Table 13: Per-task accuracies on eight zero-shot common-sense reasoning tasks.
Model Size ARC-easy ARC-challenge BoolQ PIQA SIQA HellaSwag OBQA WinoGrande Avg.
<150M
SmolLM2-135M 135M 58.9 34.5 60.5 68.9 43.5 43.6 41.1 52.8 50.5
MobileLLM-R1-140M-base 140M 47.3 32.5 57.2 62.5 42.6 32.9 31.5 51.0 44.7
150M – 400M
Gemma-3-270M-pt 268M 57.7 31.5 56.5 68.5 43.1 41.4 34.6 53.9 48.4
SmolLM2-360M 362M 67.9 42.0 63.4 72.1 44.7 56.2 47.4 59.0 56.6
MobileLLM-R1-360M-base 359M 59.8 36.0 65.9 68.3 45.0 42.7 34.4 53.5 50.7
400M – 1B
Qwen2.5-0.5B 494M 63.3 36.1 56.9 69.4 44.6 50.7 40.7 56.1 52.2
Qwen3-0.6B-Base 596M 66.2 40.3 69.4 69.7 45.5 52.9 40.7 57.9 55.3
MobileLLM-R1-950M-base 949M 66.3 45.3 79.6 74.2 48.4 57.2 39.8 58.2 58.6
>1B
Gemma-3-1B-pt 1.00B 71.6 37.7 66.3 74.8 45.4 61.9 41.9 58.3 57.2
LLaMA3.2-1B 1.24B 64.8 42.9 64.7 74.9 44.8 64.4 50.0 60.4 58.4
OLMo-2-0425-1B 1.48B 72.2 44.0 68.8 74.5 48.4 66.0 46.0 63.4 60.4
Qwen2.5-1.5B 1.54B 70.4 41.2 73.9 72.9 48.2 65.2 36.3 61.2 58.7
SmolLM2-1.7B 1.71B 73.6 49.5 72.0 77.6 46.1 70.7 47.8 64.4 62.7
Qwen3-1.7B-Base 2.03B 72.9 47.0 77.9 74.7 47.6 64.2 43.3 62.4 61.3
In smaller models, MobileLLM-R1 is not explicitly optimized for the common-sense reasoning benchmarks, which explains why SmolLM attains higher performance in this regime. In the 400M–1B parameter range, MobileLLM-R1 outperforms Qwen models and Gemma-3-1B. It is important to note that these common-sense reasoning benchmarks primarily evaluate factual, social, or physical commonsense understanding rather than tasks requiring extended chain-of-thought reasoning.

### v3

@@ B.1 Detailed Accuracy Comparison
Table 8 demonstrates the performance of sub-billion and near-billion parameter reasoning base models across six widely used benchmarks. The results demonstrate that the proposed MobileLLM-R1 models consistently outperforms prior open-source models of comparable or larger scale. In the <<150M regime, MobileLLM-R1-140M delivers substantial gains over SmolLM2-135M, particularly on GSM8K (16.3 vs. 1.8) and HumanEval (15.9 vs. 0.0). Similarly, MobileLLM-R1-360M surpasses both Gemma-3-270M-pt and SmolLM2-360M, achieving more than double the accuracy on MATH500 and GSM8K. At the higher end, MobileLLM-R1-950M matches or exceeds the performance of larger models: it improves upon Qwen3-0.6B in code generation (HumanEval: 46.3 vs. 30.5) and commonsense reasoning (58.6 vs. 55.3), while remaining competitive on math and programming tasks. These results highlight that carefully curated training data, rather than sheer scale, is the key driver of reasoning performance in lightweight models.
@@
Appendix C On-device Profiling
We benchmark the latency of MobileLLM-R1 models (140M, 360M, 950M) and LLaMA-3 models (1B, 3B, 8B) using ExecuTorch on a Samsung Galaxy S22 (8 GB RAM). In this experiment, we run each model with a warm-up phase and then average the results over three iterations using the same prompt. On-device applications usually adopt quantized settings. To reflect typical on-device deployment, models are benchmarked under quantized settings: linear layers are quantized using 8-bit dynamic activations and 4-bit weights. For the 125M model, a group size of 32 is adopted since its hidden dimension is not divisible by 128, whereas all larger models use a group size of 128. Embedding layers are quantized to 4-bit with a group size of 32.
@@ Appendix C On-device Profiling
Table 10 reports the measured generation throughput (tokens per second) of MobileLLM-R1 models and LLaMA-3 baselines across different context lengths on the Samsung Galaxy S22. The results confirm the clear efficiency advantage of smaller models: the 140M variant sustains over 100 tokens/s up to 8k tokens, while even the 360M and 950M models maintain practical throughput across medium-length contexts. In contrast, larger models such as LLaMA-3.2-3B and LLaMA-3.1-8B encounter out-of-memory failures beyond short contexts, and the LLaMA-3.2-1B model runs at less than one-quarter the speed of MobileLLM-R1-140M. Moreover, long-context settings (16k–32k) quickly exhaust memory capacity for billion-scale models, highlighting the prohibitive cost of KV-cache storage at scale. These findings demonstrate that sub-billion parameter models not only enable reasoning with reduced memory overhead but also achieve far higher generation speed, making them far better suited for on-device long-context inference.
Appendix D Ablation Analysis and Insights
D.1 Learning Rate Affects Representation Learning
@@
Therefore, we investigate structural diagnostics of pretrained representations that can act as low-cost proxies for downstream generalization. We draw inspiration from RankMe Garrido et al. (2023), originally proposed in the context of vision self-supervised finetuning:
$\text{RankMe}(Z)=\exp\left(-\sum_{k=1}^{\min(N,K)}p_{k}\log p_{k}\right),\quad\text{where }p_{k}=\frac{\sigma_{k}(Z)}{\sum_{i=1}^{\min(N,K)}\sigma_{i}(Z)}+\epsilon,$ (8)
@@
D.2 Discussion: Whether to Use RL or Not
D.3 Computational Overhead of the Data Curation Pipeline
To provide a detailed characterization of the computational overhead of our method, we measure and report the data curation costs, including representative dataset construction, influence score computation and leave-one-out ablation.
  • Hierarchical Rejection Sampling for producing representative datasets: $\sim$200 GPU hours.
  • Pre-training influence computation: $\sim$5,100 GPU hours.
  • Mid-training influence computation: $\sim$1,100 GPU hours.
  • Leave-one–out ablation: $\sim$50 GPU hours per dataset, $\sim$ 400 GPU hours in total.
  • Total: $\sim$6,800 GPU hours.
Training cost:
  • Pre-training: $\sim$28,000 GPU hours.
  • Mid-training: $\sim$10,000 GPU hours.
  • Post-training: $\sim$600 GPU hours.
  • Total: $\sim$38600 GPU hours.
The detailed training cost can be found in Table 4. Our analysis indicates that the data curation pipeline itself is also computationally efficient: the combined cost of data curation and ablation experiments constitutes only approximately 15% of the total cost of data curation and model training.
Appendix E Use of LLM
We leverage large language models (LLMs) to assist in drafting and polishing written content. Specifically, LLMs are used to improve clarity, coherence, and readability of the text, while ensuring that technical accuracy and intended meaning are preserved. All outputs generated by the LLM are carefully reviewed and edited by the authors to maintain factual correctness and align with the scientific content. The LLM serves as a tool to support writing efficiency, not to generate original research ideas or conclusions.